# Limits on planetary-mass primordial black holes from the OGLE high-cadence survey of the Magellanic Clouds

### v1

@@ 4.1 OGLE-LMC-20
The light curve can be modeled by a PSPL model with a very short Einstein timescale $t_{\rm E}=0.104^{+0.044}_{-0.021}$ day. However, its rising part is unavailable, rendering it impossible to check for the symmetry of the bump, and raising the possibility that the object was not the microlensing event at all. In fact, two lines of evidence indicate that the observed brightening was a stellar flare.
@@
The star was also detected by the VISTA survey of the Magellanic Clouds (VMC; Cioni et al. 2011; Rubele et al. 2018). Its near-infrared magnitudes ($Y=19.093\pm 0.043$, $J=18.622\pm 0.049$, $K_{\rm s}=17.731\pm 0.094$) are also consistent with those expected from late K or early M dwarfs (Pecaut & Mamajek, 2013).
Second, the proper motion of the star is inconsistent with the LMC proper motion. This object is not included in the Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2021). However, we were able to measure its proper motion using the archival OGLE data, as part of the OGLE-Uranus project (Udalski et al., in prep.). The proper motion of the star is $(\mu_{\alpha},\mu_{\delta})=(0.42\pm 0.32,3.56\pm 0.40)$ mas yr-1. On the other hand, the LMC proper motion, defined here as the mean proper motion of nearby red clump stars [^3] We used stars located in the CMD region defined by $|I-I_{\rm RC}|\leq 0.6$ and $|(V-I)-(V-I)_{\rm RC}|\leq 0.2$, where $I_{\rm RC}$ and $(V-I)_{\rm RC}$ are the apparent magnitude and color of the red clump centroid. is $(\mu_{\alpha},\mu_{\delta})_{\rm RC}=(1.82\pm 0.35,0.45\pm 0.44)$ mas yr-1. (The latter uncertainties quantify the rms scatter around the mean.) The proper motion of the star in the East and North direction differs by $3.0\sigma$ and $5.2\sigma$, respectively, from the LMC proper motion. The $p$-value of the hypothesis that the proper motion is consistent with that of the LMC is just $5\times 10^{-10}$, so this hypothesis can be ruled out at 6.2-sigma confidence level.
@@ 5 Event Detection Efficiency
We carried out extensive light-curve level simulations to calculate the event detection efficiency in the high-cadence survey. Their methodology is similar to that used in our previous works (Mróz et al., 2019, 2024a). It involves injecting the microlensing signal on top of the light curves of stars randomly drawn from the database, thereby preserving sampling, variability, and noise in the original data that would otherwise be difficult to simulate. We then check if the simulated light curves pass all selection criteria discussed in Section 3 (and summarized in Table 2).
Because we need to take into account finite-source effects, the detection efficiency $\varepsilon(t_{\rm E},\rho)$ is calculated as a function of both the Einstein timescale $t_{\rm E}$ and the normalized source radius $\rho$. Finite-source effects affect the detectability of microlensing events in at least three ways. First, if $\rho\gg 1$, the duration of the event is set by the time needed for the lens to cross the source’s surface, making it easier to detect small $\theta_{\rm E}$ (short $t_{\rm E}$) events. At the same time, the amplitude of the event (in magnitudes) decreases with increasing source radius as $2.171/\rho^{2}$. If $\rho>6.6$, the amplitude of the event falls below 0.05 mag, which was a threshold adopted for selecting candidate events (Table 2). Finally, the cross-section to microlensing is no longer set by the angular Einstein radius but rather the angular size of the source star $\theta_{*}$.
The simulations were carried out for nine values of the normalized source radius ($\log\rho\in\{-3,-2,-1,0,0.2,0.4,0.6,0.8,1\}$). The peak time $t_{0}$ was drawn from a uniform distribution from the range $4502,459,750\leq t_{0}\leq 2,460,$ 450 (i.e., between 2022 Jun 19.5 and 2024 May 19.5). The impact parameter was drawn from a range $0\leq u_{0}\leq u_{\rm max}$, where $u_{\rm max}=\max\{1,2\rho\}$ to take into account all possible source-crossing events; and timescales were drawn from a log-uniform distribution from the range $10^{-2}$ to $10^{2}$ day. The dimensionless blending parameter was drawn from empirical distributions created by matching stars detected in OGLE reference frames and high-resolution Hubble Space Telescope images (c.f., Mróz et al. 2024a).
@@ 6 Limits on Planetary-Mass PBHs
We calculate the limits on the abundance of planetary-mass PBHs in dark matter as a function of their mass $M$, assuming a monochromatic (delta-function) mass function. We denote the fraction of dark matter in the form of PBHs as $f=M_{\rm PBH}/M_{\rm DM}$, where $M_{\rm DM}$ is the total mass of dark matter halo, and $M_{\rm PBH}$ is the total mass of PBHs within it. Our methodology closely follows that presented in Mróz et al. (2024b).
We use the Milky Way and LMC disk and halo models to calculate the differential event rate $d^{2}\Gamma/dt_{\rm E}d\rho$, and calculate the expected number of events
@@
where $N_{\rm s}$ is the number of source stars brighter than $I=22$, and $\Delta t=700$ day is the duration of the survey. The number of source stars used in calculations is reported in Table 1, and taken from Mróz et al. (2024a) for the LMC and Mróz et al. (2025, in prep.) for the SMC. The integral is evaluated by generating an ensemble of $10^{6}$ events using the method described by Clanton & Gaudi (2014). The finite-source effects are taken into account by assigning each simulated event a random source brightness $I_{\rm s}$ drawn from the luminosity function of a given field. Then, the angular radius of the source is calculated from
@@
where $I_{\rm RC}$ is the mean brightness of red clump stars in this field, and $\rho=\theta_{*}/\theta_{\rm E}$. The angular radius of red clump stars is calculated using data from Nataf et al. (2021).
The number of microlensing events expected to be detected in our survey, assuming that the entire dark matter is made of compact objects, is presented in Figure 7. We sum the contributions from the Milky Way halo (which is assumed to follow the Cautun et al. (2020) contracted halo model with the total mass of $0.97\times 10^{12}\,M_{\odot}$ within 200 kpc) and the LMC halo (which is modeled by a Hernquist profile with a total mass of $1.49\times 10^{11}\,M_{\odot}$; Erkal et al. 2019). These models are described in more detail in Mróz et al. (2024b). We neglect the contribution from the SMC dark matter halo because of its small mass ($\approx 6.5\times 10^{9}\,M_{\odot}$; Bekki & Stanimirović 2009), relatively small number of sources in the SMC (four times smaller than in the LMC), and shorter duration of the SMC high-cadence survey (236 vs. 435 nights).
@@
On the other hand, if we take into account only known stellar populations in the Milky Way disk, we should have detected from 0.4 to 0.9 microlensing events depending on the adopted disk model (Han & Gould (2003) or Cautun et al. (2020), respectively). The expected number of self-lensing events in the LMC is only 0.1. Thus, the total number of expected events is 0.5–1.0, in good agreement with one event detected. The timescale of this event ($t_{\rm E}=90^{+13}_{-10}$ day) also matches those expected from the Milky Way and LMC stellar populations (see Extended Data Figure 6 in Mróz et al. 2024b).
Assuming that OGLE-LMC-20 originates from the lens in the Milky Way disk or in the LMC, we calculate the posterior distribution for the frequency of PBHs in dark matter $f$ from the Bayes’ formula
@@
is the likelihood function, and $P_{0}(f)$ is a flat (uniform) prior on $f$. This calculation is repeated for 121 logarithmically spaced masses ranging from $10^{-10}\,M_{\odot}$ to $10^{2}\,M_{\odot}$ using the methodology developed by Mróz et al. (2024b).

### v2

@@ 4.1 OGLE-LMC-20
The light curve can be modeled by a PSPL model with a very short Einstein timescale $t_{\rm E}=0.104^{+0.044}_{-0.021}$ day. However, its rising part is unavailable, rendering it impossible to check for the symmetry of the bump and raising the possibility that the object was not the microlensing event at all. In fact, two lines of evidence indicate that the observed brightening was a stellar flare.
@@
The star was also detected by the VISTA survey of the Magellanic Clouds (Cioni et al. 2011; Rubele et al. 2018). Its near-infrared magnitudes ($Y=19.093\pm 0.043$, $J=18.622\pm 0.049$, $K_{\rm s}=17.731\pm 0.094$) are also consistent with those expected from late K or early M dwarfs (Pecaut & Mamajek, 2013).
Second, the proper motion of the star is inconsistent with the LMC proper motion, as shown in Figure 5. This object is not included in the Gaia Data Release 3 (Gaia Collaboration et al., 2016, 2021). However, we were able to measure its proper motion using the archival OGLE data as part of the OGLE-Uranus project (Udalski et al., in preparation). The proper motion of the star is $(\mu_{\alpha},\mu_{\delta})=(0.42\pm 0.32,3.56\pm 0.40)$ mas yr-1. On the other hand, the LMC proper motion, defined here as the mean proper motion of nearby red clump stars, [^3] We used stars located in the CMD region defined by $|I-I_{\rm RC}|\leq 0.6$ and $|(V-I)-(V-I)_{\rm RC}|\leq 0.2$, where $I_{\rm RC}$ and $(V-I)_{\rm RC}$ are the apparent magnitude and color of the red clump centroid. is $(\mu_{\alpha},\mu_{\delta})_{\rm RC}=(1.82\pm 0.35,0.45\pm 0.44)$ mas yr-1. (The latter uncertainties quantify the rms scatter around the mean.) The proper motion of the star in the east and north direction differs by $3.0\sigma$ and $5.2\sigma$, respectively, from the LMC proper motion. The $p$-value of the hypothesis that the proper motion is consistent with that of the LMC is just $5\times 10^{-10}$, so this hypothesis can be ruled out at $6.2\sigma$ confidence level.
@@ 5 Event Detection Efficiency
We carried out extensive light-curve level simulations to calculate the event detection efficiency in the high-cadence survey. Their methodology is similar to that used in our previous works (Mróz et al., 2019, 2024a). It involves injecting the microlensing signal on top of the light curves of stars randomly drawn from the database, thereby preserving sampling, variability, and noise in the original data that would otherwise be difficult to simulate. We then check if the simulated light curves pass all the selection criteria discussed in Section 3 (and summarized in Table 2).
Because we need to take into account finite-source effects, the detection efficiency $\varepsilon(t_{\rm E},\rho)$ is calculated as a function of both the Einstein timescale $t_{\rm E}$ and the normalized source radius $\rho$. Finite-source effects affect the detectability of microlensing events in at least three ways. First, if $\rho\gg 1$, the duration of the event is set by the time needed for the lens to cross the source’s surface, making it easier to detect small $\theta_{\rm E}$ (short $t_{\rm E}$) events. At the same time, the amplitude of the event (in magnitudes) decreases with increasing source radius as $2.171/\rho^{2}$. If $\rho>6.6$, the amplitude of the event falls below 0.05 mag, which was a threshold adopted for selecting candidate events (Table 2). Finally, the cross section to microlensing is no longer set by the angular Einstein radius but rather the angular size of the source star $\theta_{*}$.
The simulations were carried out for nine values of the normalized source radius ($\log\rho\in\{-3,-2,-1,0,0.2,0.4,0.6,0.8,1\}$). The peak time $t_{0}$ was drawn from a uniform distribution from the range $4502,459,750\leq t_{0}\leq 2,460,$ 450 (i.e., between 2022 June 19.5 and 2024 May 19.5). The impact parameter was drawn from a range $0\leq u_{0}\leq u_{\rm max}$, where $u_{\rm max}=\max\{1,2\rho\}$ to take into account all possible source-crossing events, and timescales were drawn from a log-uniform distribution from the range $10^{-2}$ to $10^{2}$ days. The dimensionless blending parameter was drawn from empirical distributions created by matching stars detected in OGLE reference frames and high-resolution Hubble Space Telescope images (see, Mróz et al. 2024a).
@@ 6 Limits on Planetary-mass PBHs
We calculate the limits on the abundance of planetary-mass PBHs in dark matter as a function of their mass $M$, assuming a monochromatic (delta-function) mass function. We denote the fraction of dark matter in the form of PBHs as $f=M_{\rm PBH}/M_{\rm DM}$, where $M_{\rm DM}$ is the total mass of dark matter halo and $M_{\rm PBH}$ is the total mass of PBHs within it. Our methodology closely follows that presented in Mróz et al. (2024b).
We use the Milky Way and LMC disk and halo models to calculate the differential event rate $d^{2}\Gamma/dt_{\rm E}d\rho$ and calculate the expected number of events:
@@
where $N_{\rm s}$ is the number of source stars brighter than $I=22$ and $\Delta t=700$ day is the duration of the survey. The number of source stars used in calculations is reported in Table 1 and taken from Mróz et al. (2024a) for the LMC and Mróz et al. (2025, in preparation) for the SMC. The integral is evaluated by generating an ensemble of $10^{6}$ events using the method described by Clanton & Gaudi (2014). The finite-source effects are taken into account by assigning each simulated event a random source brightness $I_{\rm s}$ drawn from the luminosity function of a given field. Then, the angular radius of the source is calculated from
@@
where $I_{\rm RC}$ is the mean brightness of red clump stars in this field and $\rho=\theta_{*}/\theta_{\rm E}$. The angular radius of red clump stars is calculated using data from Nataf et al. (2021).
The number of microlensing events expected to be detected in our survey, assuming that the entire dark matter is made of compact objects, is presented in Figure 7. We sum the contributions from the Milky Way halo (which is assumed to follow the Cautun et al. (2020) contracted halo model with the total mass of $0.97\times 10^{12}\,M_{\odot}$ within 200 kpc) and the LMC halo (which is modeled by a Hernquist profile with a total mass of $1.49\times 10^{11}\,M_{\odot}$; Erkal et al. 2019). These models are described in more detail in Mróz et al. (2024b). We neglect the contribution from the SMC dark matter halo because of its small mass ($\approx 6.5\times 10^{9}\,M_{\odot}$; Bekki & Stanimirović 2009), relatively small number of sources in the SMC (four times smaller than in the LMC), and shorter duration of the SMC high-cadence survey (236 versus 435 nights).
@@
On the other hand, if we take into account only known stellar populations in the Milky Way disk, we should have detected from 0.4 to 0.9 microlensing events depending on the adopted disk model (Han & Gould 2003 or Cautun et al. 2020, respectively). The expected number of self-lensing events in the LMC is only 0.1. Thus, the total number of expected events is 0.5–1.0, in good agreement with one event detected. The timescale of this event ($t_{\rm E}=90^{+13}_{-10}$ day) also matches that expected from the Milky Way and LMC stellar populations (see Extended Data Figure 6 in Mróz et al. 2024b).
Assuming that OGLE-LMC-20 originates from the lens in the Milky Way disk or in the LMC, we calculate the posterior distribution for the frequency of PBHs in dark matter $f$ from the Bayes’ formula:
@@
is the likelihood function and $P_{0}(f)$ is a flat (uniform) prior on $f$. This calculation is repeated for 121 logarithmically spaced masses ranging from $10^{-10}\,M_{\odot}$ to $10^{2}\,M_{\odot}$ using the methodology developed by Mróz et al. (2024b).